OpenStudy (anonymous):

Can the sum of two polygons in x, both of degree 4, be of degree 3? If so give an example.

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@Hero @ganeshie8

OpenStudy (anonymous):

Yes !

OpenStudy (anonymous):

example??

OpenStudy (anonymous):

Let \[P(x)=x^4+3x^3+5x^2-1\text{ and } Q(x)=-x^4+2x^2+17\] We have : \[\deg P=\deg Q=4\] But : \[\deg(P+Q)=3\]

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

so can i use that as my example do you mind?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

you mind?